The length of a rectangle is 8 units greater than the width, x. Write an expression giving the area, A, of the rectangle.
Answer by cleomenius(959)

width = x
length = x+ 8
(x)(x + 8) = Area
x^2 + 8x = Area.
Cleomenius.